# In this paper, we show that the constant-dimensional Weisfeiler–Leman algorithm for groups (Brachter & Schweitzer, LICS 2020) can be fruitfully used to improve parallel complexity upper bounds on isomorphism testing for several families of groups. In particular, we show:

### v6

@@ 1 Introduction
[…]
@@
Let $k\geq 2$ and $r\geq 0$ be integers, and let $J\in\{I,II\}$. The $(k,r)$-WL Version $J$ algorithm for groups is obtained by running $k$-WL Version $J$ for $r$ rounds. Here, the initial coloring counts as the first round ($r=1$). By conventional definition, each of the WL versions distinguish two groups in 00 rounds ($r=0$) iff the groups do not have the same order.
@@
Let $G=G_{1}\times\cdots\times G_{d}$ be a decomposition into indecomposable direct factors, let $k\geq 5$, and let $r:=r(n)$. If $G$ and $H$ are not distinguished by $(k,r+O(\log n))$-WL Version II, then there exist direct factors $H_{i}\leq H$ such that $H=H_{1}\times\cdots\times H_{d}$ such that for all $i\in[d]$, $G_{i}$ and $H_{i}$ are not distinguished by $(k-1,r)$-WL Version II.
@@
[…]
@@
Here, $\beta_{1}\textsf{MAC}^{0}(\textsf{FOLL})$ denotes the composition of an FOLL circuit with a $\beta_{1}\textsf{MAC}^{0}$ circuit. We refer to Section 2.5 for formal details of these complexity classes. Here, we will briefly sketch the algorithm in hopes of providing some intuition regarding the complexity bounds. We may easily check in $\textsf{AC}^{0}$ whether a group is Abelian. So suppose that $G,H$ are Abelian groups. We first compute the order of each element, for both groups by an $\mathsf{AC}$ circuits of depth $O(\log\log n)$, that is, in the class $\mathsf{FOLL}$ [BKLM01]. If $G\not\cong H$, then there exists $d\in\mathbb{N}$ where, without loss of generality, $G$ has more elements of order $d$ than $H$. We may use $O(\log n)$ non-deterministic bits to guess such a $d$ (this is the $\beta_{1}$), and a single Majority gate to compare the number of elements of order $d$ in $G$ vs. $H$ (this is the $\textsf{MAC}^{0}$ circuit). We refer to the proof (see Theorem 7.12) for full details.
@@ Methods.
Lemma 1.13 (Rank lemma).
@@
One application of our Rank Lemma is that WL identifies verbal subgroups where the words are easily identified. Given a set of words $w_{1}(x_{1},\dotsc,x_{n}),\dotsc,w_{m}(x_{1},\ldots,x_{m})$, the corresponding verbal subgroup is the subgroup generated by $\{w_{i}(g_{1},\dotsc,g_{n}):i=1,\dotsc,m,g_{j}\in G\}$. One example that we use in our results is the commutator subgroup. If Duplicator chooses a bijection $f:G\to H$ such that $f([x,y])$ is not a commutator in $H$, then Spoiler pebbles $[x,y]\mapsto f([x,y])$ and wins in two additional rounds. Thus, by our Rank Lemma, if Duplicator does not map the commutator subgroup $[G,G]$ to the commutator subgroup $[H,H]$, then Duplicator wins with $1$ additional pebble and $O(\log n)$ additional rounds.
@@
Combinatorial techniques, such as individualization with Weisfeiler–Leman refinement, have also been incredibly successful in GI, yielding efficient isomorphism tests for several families [GV06, KPS19, GK21, GK19, GN21, BW13, CST13]. Weisfeiler–Leman is also a key subroutine in Babai’s quasipolynomial-time isomorphism test [Bab16]. Despite the successes of such combinatorial techniques, they are known to be insufficient to place GI into P [CFI92, NS18]. In contrast, the use of combinatorial techinques for GpI is relatively new [LQ17, BGL+19, BS20, BS22], and it is a central open problem as to whether such techniques are sufficient to improve even the long-standing upper-bound of $n^{\Theta(\log n)}$ runtime.
@@ 2.1 Groups
We say that a normal subgroup $N\trianglelefteq G$ splits in $G$ if there exists a subgroup $H\leq G$ such that $H\cap N=\{1\}$ and $G=HN$. The conjugation action of $H$ on $N$ allows us to express multiplication of $G$ in terms of pairs $(h,n)\in H\times N$. We note that the conjugation action of $H$ on $N$ induces a group homomorphism $\theta:H\to\operatorname{Aut}(N)$ mapping $h\mapsto\theta_{h}$, where $\theta_{h}:N\to N$ sends $\theta_{h}(n)=hnh^{-1}$. So given $(H,N,\theta)$, we may define the group $H\ltimes_{\theta}N$ on the set $\{(h,n):h\in H,n\in N\}$ with the product $(h_{1},n_{1})(h_{2},n_{2})=(h_{1}h_{2},\theta_{h_{2}^{-1}}(n_{1})n_{2})$. We refer to the decomposition $G=H\ltimes_{\theta}N$ as a semidirect product demoposition. When the action $\theta$ is understood, we simply write $G=H\ltimes N$.
@@ 2.2 Weisfeiler–Leman
We begin by recalling the Weisfeiler–Leman algorithm for graphs, which computes an isomorphism-invariant coloring. Let $\Gamma$ be a graph, and let $k\geq 2$ be an integer. The $k$-dimensional Weisfeiler–Leman, or $k$-WL, algorithm begins by constructing an initial coloring $\chi_{k,1}:V(\Gamma)^{k}\to\mathcal{K}$, where $\mathcal{K}$ is our set of colors, by assigning each $k$-tuple a color based on its isomorphism type. That is, two $k$-tuples $(v_{1},\ldots,v_{k})$ and $(u_{1},\ldots,u_{k})$ receive the same color under $\chi_{k,0}$ iff the map $v_{i}\mapsto u_{i}$ (for all $i\in[k]$) is an isomorphism of the induced subgraphs $\Gamma[\{v_{1},\ldots,v_{k}\}]$ and $\Gamma[\{u_{1},\ldots,u_{k}\}]$ and for all $i,j$, $v_{i}=v_{j}\Leftrightarrow u_{i}=u_{j}$.
@@
The coloring computed at the $(r+1)$st iteration, denoted $\chi_{k,r+1}$, stores the color of the given $k$-tuple $\overline{v}$ at the $r$th iteration, as well as the colors under $\chi_{k,r}$ of the $k$-tuples obtained by substituting a single vertex in $\overline{v}$ for another vertex $x$. We examine this multiset of colors over all such vertices $x$. This is formalized as follows:
@@
The count-free variant of $k$-WL works identically as the classical variant, except at the refinement step, we consider the set of colors rather than the full multi-set. Precisely:
@@
Similarly, the count-free variant of $(k,r)$-WL distinguishes $G$ from the graph $H$ if there exists a color class $C$ and some $\bar{x}\in V(G)^{k}$ where $\chi_{k,r}(\bar{x})=C$, but for all $\bar{y}\in V(H)^{k}$, $\chi_{k,r}(\bar{y})\neq C$. We say that the classical counting (resp., count-free) $(k,r)$-WL identifies $G$ if for all $H\not\cong G$, $(k,r)$-WL (resp., count-free $(k,r)$-WL) distinguishes $G$ from $H$.
@@ 2.3 Pebbling Game
We recall the bijective pebble game introduced by Hella [Hel89, Hel96] for WL on graphs. This game is often used to show that two graphs $X$ and $Y$ cannot be distinguished by $k$-WL. The game is an Ehrenfeucht–Fraïssé game (c.f., [EFT94, Lib04]), with two players: Spoiler and Duplicator. We begin with $k+1$ pairs of pebbles. Prior to the start of the game, each pebble pair $(p_{i},p_{i}^{\prime})$ is initially placed either beside the graphs or on a given pair of vertices $v_{i}\mapsto v_{i}^{\prime}$ (where $v_{i}\in V(X),v_{i}^{\prime}\in V(Y)$). Each round $r$ proceeds as follows.
@@
Let $v_{1},\ldots,v_{m}$ be the vertices of $X$ pebbled at the end of step 1 at round $r$ of the game, and let $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ be the corresponding pebbled vertices of $Y$. Spoiler wins precisely if the map $v_{\ell}\mapsto v_{\ell}^{\prime}$ does not extend to an isomorphism of the induced subgraphs $X[\{v_{1},\ldots,v_{m}\}]$ and $Y[\{v_{1}^{\prime},\ldots,v_{m}^{\prime}\}]$. Duplicator wins otherwise. Spoiler wins, by definition, at round 00 if $X$ and $Y$ do not have the same number of vertices. We note that $\overline{v}\in X^{k}$ and $\overline{v^{\prime}}\in Y^{k}$ are not distinguished by the first $r$ rounds of $k$-WL if and only if Duplicator wins the first $r$ rounds of the $(k+1)$-pebble game starting from the configuration $\overline{v}\mapsto\overline{v^{\prime}}$ [Hel89, Hel96, CFI92].
@@ 2.5 Complexity Classes
[…]
The complexity class $\textsf{MAC}^{0}$ is the set of languages decidable by constant-depth uniform circuit familes with a polynomial number of AND, OR, and NOT gates, and at most one Majority gate. The class $\textsf{MAC}^{0}$ was introduced (but not so named) in [ABFR91], where it was shown that $\textsf{MAC}^{0}\subsetneq\textsf{TC}^{0}$. This class was subsequently given the name $\textsf{MAC}^{0}$ in [JKS02].
@@ 3.1 Additional preliminaries for groups with Abelian normal Hall subgroup
A $\mathbb{Z}H$-module is an abelian group $N$ together with an action of $H$ on $N$, given by a group homomorphism $\theta\colon H\to\operatorname{Aut}(N)$. We refer to these as “$H$-modules.” A submodule of an $H$-module $N$ is a subgroup $N^{\prime}\leq N$ such that the action of $H$ on $N^{\prime}$ sends $N^{\prime}$ into itself, and thus the restriction of the action of $H$ to $N^{\prime}$ gives $N^{\prime}$ the structure of an $H$-module compatible with that on $N$. Given a subset $S\subseteq N$, the smallest $H$-submodule containing $S$ is denoted $\langle S\rangle_{H}$, and is referred to as the $H$-submodule generated by $S$. An $H$-module generated by a single element is called cyclic. Note that a cyclic $H$-module $N$ need not be a cyclic Abelian group.
Two $H$-modules $N,N^{\prime}$ are isomorphic (as $H$-modules), denoted $N\cong_{H}N^{\prime}$, if there is a group isomorphism $\varphi\colon N\to N^{\prime}$ that is $H$-equivariant, in the sense that $\varphi(\theta(h)(n))=\theta^{\prime}(h)(\varphi(n))$ for all $h\in H,n\in N$. An $H$-module $N$ is decomposable if $N\cong_{H}N_{1}\oplus N_{2}$ where $N_{1},N_{2}$ are nonzero $H$-modules (and the direct sum can be thought of as a direct sum of Abelian groups); otherwise $N$ is indecomposable. An equivalent characterization of $N$ being decomposable is that there are nonzero $H$-submodules $N_{1},N_{2}$ such that $N=N_{1}\oplus N_{2}$ as Abelian groups (that is, $N$ is generated as a group by $N_{1}$ and $N_{2}$, and $N_{1}\cap N_{2}=0$). The Remak–Krull–Schmidt Theorem says that every $H$-module decomposes as a direct sum of indecomposable modules, and that the multiset of $H$-module isomorphism types of the indecomposable modules appearing is independent of the choice of decomposition, that is, it depends only on the $H$-module isomorphism type of $N$. We may thus write

### v7

@@ 1 Introduction
The Group Isomorphism problem (GpI) takes as input two finite groups $G$ and $H$, and asks if there exists an isomorphism $\varphi:G\to H$. When the groups are given by their multiplication (a.k.a. Cayley) tables, it is known that GpI belongs to $\textsf{NP}\cap\textsf{coAM}$. The generator-enumerator algorithm, attributed to Tarjan in 1978 [Mil78], has time complexity $n^{\log_{p}(n)+O(1)}$, where $n$ is the order of the group and $p$ is the smallest prime dividing $n$. In more than 40 years, this bound has escaped largely unscathed: Rosenbaum [Ros13] (see [LGR16, Sec. 2.2]) improved this to $n^{(1/4)\log_{p}(n)+O(1)}$. And even the impressive body of work on practical algorithms for this problem, led by Eick, Holt, Leedham-Green and O’Brien (e. g., [BEO02, ELGO02, BE99, CH03]) still results in an $n^{\Theta(\log n)}$-time algorithm in the general case (see [Wil19, Page 2]). In the past several years, there have been significant advances on algorithms with worst-case guarantees on the serial runtime for special cases of this problem including Abelian groups [Kav07, Vik96, Sav80], direct product decompositions [Wil12, KN09], groups with no Abelian normal subgroups [BCGQ11, BCQ12], coprime and tame group extensions [LG09, QST11, BQ12, GQ15], low-genus $p$-groups and their quotients [LW12, BMW17], Hamiltonian groups [DS19], and groups of almost all orders [DW22].
@@
Let $k\geq 2$ and $r\geq 0$ be integers, and let $J\in\{I,II\}$. The $(k,r)$-WL Version $J$ algorithm for groups is obtained by running $k$-WL Version $J$ for $r$ rounds. Here, the initial coloring counts as the first round ($r=0$). By conventional definition, each of the WL versions distinguish two group at $r=0$ if and only if the groups do not have the same order.
@@
Let $G=G_{1}\times\cdots\times G_{d}$ be a decomposition into indecomposable direct factors, let $k\geq 5$, and let $r:=r(n)$. If $G$ and $H$ are not distinguished by $(k,r+O(\log n))$-WL Version II, then there exist direct factors $H_{i}\leq H$ such that $H=H_{1}\times\cdots\times H_{d}$ such that for all $i\in[d]$, $G_{i}$ and $H_{i}$ are not distinguished by (k1,r)(k-1,r)-WL Version II.
@@
We next consider groups without Abelian normal subgroups, one of the few other classes of groups of structural importance for which a highly non-trivial polynomial-time algorithm is known. Using the individualize-and-refine paradigm, we obtain a new upper bound for not only deciding isomorphisms, but also listing isomorphisms. To state this result, recall that an SAC circuit is a Boolean circuit using $\textsf{AND},\textsf{OR}$, and NOT gates, in which the AND gates have fan-in $2$, and the OR gates have unbounded fan-in (see Section 2.5 for more details). We show that isomorphisms of groups without Abelian normal subgroups can be listed by an $\mathsf{SAC}$ circuit of $O(\log n)$ depth and $n^{O(\log\log n)}$ size. Although this does not improve upon the upper bound of P for isomorphism testing in this class of groups [BCQ12], this does parallelize the previous bound of $n^{\Theta(\log\log n)}$ runtime for listing isomorphisms [BCGQ11]. We note that for listing isomorphisms, our size bound is essentially optimal, as Babai et al. (ibid.) observed that such groups can have $n^{\Theta(\log\log n)}$ many isomorphisms. Compared to what is known for general groups, general group isomorphism can be solved by $O(\log n)$-depth $n^{O(\log n)}$-size $\mathsf{SAC}$ circuits [CTW13], and for groups without Abelian normal subgroups our result improves the size much closer to polynomial while maintaining the depth.
@@
Here, $\beta_{1}\textsf{MAC}^{0}(\textsf{FOLL})$ denotes the composition of an FOLL circuit with a $\beta_{1}\textsf{MAC}^{0}$ circuit. We refer to Section 2.5 for formal details of these complexity classes. Here, we will briefly sketch the algorithm in hopes of providing some intuition regarding the complexity bounds. We may easily check in $\textsf{AC}^{0}$ whether a group is Abelian. So suppose that $G,H$ are Abelian groups. We first compute the order of each element, for both groups by an $\mathsf{AC}$ circuit of depth $O(\log\log n)$, that is, in the class $\mathsf{FOLL}$ [BKLM01]. If $G\not\cong H$, then there exists $d\in\mathbb{N}$ where, without loss of generality, $G$ has more elements of order $d$ than $H$. We may use $O(\log n)$ non-deterministic bits to guess such a $d$ (this is the $\beta_{1}$), and a single Majority gate to compare the number of elements of order $d$ in $G$ vs. $H$ (this is the $\textsf{MAC}^{0}$ circuit). We refer to the proof (see Theorem 7.12) for full details.
@@ Methods.
Lemma 1.13 (Rank lemma; cf. Lemma 4.3).
@@
One application of our Rank Lemma is that WL identifies verbal subgroups where the words are easily identified. Given a set of words $w_{1}(x_{1},\dotsc,x_{n}),\dotsc,w_{m}(x_{1},\ldots,x_{n})$, the corresponding verbal subgroup is the subgroup generated by $\{w_{i}(g_{1},\dotsc,g_{n}):i=1,\dotsc,m,g_{j}\in G\}$. One example that we use in our results is the commutator subgroup. If Duplicator chooses a bijection $f:G\to H$ such that $f([x,y])$ is not a commutator in $H$, then Spoiler pebbles $x,y]\mapsto f([x,y])$ and wins in two additional rounds. Thus, by our Rank Lemma, if Duplicator does not map the commutator subgroup $[G,G]$ to the commutator subgroup $[H,H]$, then Duplicator wins with $1$ additional pebble and $O(\log n)$ additional rounds.
@@
Combinatorial techniques, such as individualization with Weisfeiler–Leman refinement, have also been incredibly successful in GI, yielding efficient isomorphism tests for several families [GV06, KPS19, GK21, GK19, GN23, BW13, CST13]. Weisfeiler–Leman is also a key subroutine in Babai’s quasipolynomial-time isomorphism test [Bab16]. Despite the successes of such combinatorial techniques, they are known to be insufficient to place GI into P [CFI92, NS18]. In contrast, the use of combinatorial techniques for GpI is relatively new [LQ17, BGL+19, BS20, BS22], and it is a central open problem as to whether such techniques are sufficient to improve even the long-standing upper-bound of $n^{\Theta(\log n)}$ runtime.
@@ 2.1 Groups
We say that a normal subgroup $N\trianglelefteq G$ splits in $G$ if there exists a subgroup $H\leq G$ such that $H\cap N=\{1\}$ and $G=HN$. The conjugation action of $H$ on $N$ allows us to express multiplication of $G$ in terms of pairs $(h,n)\in H\times N$. We note that the conjugation action of $H$ on $N$ induces a group homomorphism $\theta:H\to\operatorname{Aut}(N)$ mapping $h\mapsto\theta_{h}$, where $\theta_{h}:N\to N$ sends $\theta_{h}(n)=hnh^{-1}$. So given $(H,N,\theta)$, we may define the group $H\ltimes_{\theta}N$ on the set $\{(h,n):h\in H,n\in N\}$ with the product $(h_{1},n_{1})(h_{2},n_{2})=(h_{1}h_{2},\theta_{h_{2}^{-1}}(n_{1})n_{2})$. We refer to the decomposition $G=H\ltimes_{\theta}N$ as a semidirect product decomposition. When the action $\theta$ is understood, we simply write $G=H\ltimes N$.
@@ 2.2 Weisfeiler–Leman
We begin by recalling the Weisfeiler–Leman algorithm for graphs, which computes an isomorphism-invariant coloring. Let $\Gamma$ be a graph, and let $k\geq 2$ be an integer. The $k$-dimensional Weisfeiler–Leman, or $k$-WL, algorithm begins by constructing an initial coloring $\chi_{k,0}:V(\Gamma)^{k}\to\mathcal{K}$, where $\mathcal{K}$ is our set of colors, by assigning each $k$-tuple a color based on its isomorphism type. That is, two $k$-tuples $(v_{1},\ldots,v_{k})$ and $(u_{1},\ldots,u_{k})$ receive the same color under $\chi_{k,0}$ iff the map $v_{i}\mapsto u_{i}$ (for all $i\in[k]$) is an isomorphism of the induced subgraphs $\Gamma[\{v_{1},\ldots,v_{k}\}]$ and $\Gamma[\{u_{1},\ldots,u_{k}\}]$ and for all $i,j$, $v_{i}=v_{j}\Leftrightarrow u_{i}=u_{j}$.
@@
The coloring computed at the (r+1)(r+1)st iteration, denoted $\chi_{k,r+1}$, stores the color of the given $k$-tuple $\overline{v}$ at the $r$th iteration, as well as the colors under $\chi_{k,r}$ of the $k$-tuples obtained by substituting a single vertex in $\overline{v}$ for another vertex $x$. We examine this multiset of colors over all such vertices $x$. This is formalized as follows:
@@
The count-free variant of $k$-WL works identically as the classical variant, except at the refinement step, we consider the set of colors rather than the full multi-set. We re-use the notation $\chi_{k,r}$ to denote the coloring computed by count-free $(k,r)$-WL; context should make it clear whether $\chi_{k,r}$ refers to count-free or counting WL (we never use $\chi_{k,r}$ to denote the count-free coloring when discussing counting WL, nor vice versa). Precisely:
@@
Similarly, the count-free variant of $(k,r)$-WL distinguishes $G$ from the graph $H$ if $|V(G)|\neq|V(H)|$ or there exists a color class $C$ and some $\bar{x}\in V(G)^{k}$ where $\chi_{k,r}(\bar{x})=C$, but for all $\bar{y}\in V(H)^{k}$, $\chi_{k,r}(\bar{y})\neq C$. We say that the classical counting (resp., count-free) $(k,r)$-WL identifies $G$ if for all $H\not\cong G$, $(k,r)$-WL (resp., count-free $(k,r)$-WL) distinguishes $G$ from $H$. The terms distinguish and identify also extend in the natural way to WL on groups.
@@ 2.3 Pebbling Game
We recall the bijective pebble game introduced by Hella [Hel89, Hel96] for WL on graphs. This game is often used to show that two graphs $X$ and $Y$ cannot be distinguished by $k$-WL. The game is an Ehrenfeucht–Fraïssé game (cf., [EFT94, Lib04]), with two players: Spoiler and Duplicator. We begin with $k+1$ pairs of pebbles. Prior to the start of the game, each pebble pair $(p_{i},p_{i}^{\prime})$ is initially placed either beside the graphs or on a given pair of vertices $v_{i}\mapsto v_{i}^{\prime}$ (where $v_{i}\in V(X),v_{i}^{\prime}\in V(Y)$). Each round $r$ proceeds as follows.
@@
Let $v_{1},\ldots,v_{m}$ be the vertices of $X$ pebbled at the end of step 1 at round $r$ of the game, and let $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ be the corresponding pebbled vertices of $Y$. Spoiler wins precisely if the map $v_{\ell}\mapsto v_{\ell}^{\prime}$ does not extend to an isomorphism of the induced subgraphs $X[\{v_{1},\ldots,v_{m}\}]$ and $Y[\{v_{1}^{\prime},\ldots,v_{m}^{\prime}\}]$. Duplicator wins otherwise. Spoiler wins, by definition, at round 0 if $X$ and $Y$ do not have the same number of vertices. We note that $\overline{v}\in X^{k}$ and $\overline{v^{\prime}}\in Y^{k}$ are not distinguished by the first $r$ rounds of $k$-WL if and only if Duplicator wins the first $r$ rounds of the $(k+1)$-pebble game starting from the configuration $\overline{v}\mapsto\overline{v^{\prime}}$ [Hel89, Hel96, CFI92].
@@ 2.5 Complexity Classes
We assume familiarity with the complexity classes $\textsf{P},\textsf{NP},\textsf{L}$, NL, $\textsf{NC}^{k},\textsf{AC}^{k}$, and $\textsf{TC}^{k}$- we defer the reader to standard references [Zoo, AB09]. The complexity class $\textsf{SAC}^{k}$ is defined analogously to $\textsf{AC}^{k}$, except that the AND gates have bounded fan-in (while the OR gates may still have unbounded fan-in). The complexity class FOLL is the set of languages decidable by uniform circuit families with AND, OR, and NOT gates of depth $O(\log\log n)$, polynomial size, and unbounded fan-in. It is known that $\textsf{AC}^{0}\subsetneq\textsf{FOLL}\subsetneq\textsf{AC}^{1}$, and it is open as to whether FOLL is contained in NL [BKLM01].
The complexity class $\textsf{MAC}^{0}$ is the set of languages decidable by constant-depth uniform circuit families with a polynomial number of AND, OR, and NOT gates, and at most one Majority gate. The class $\textsf{MAC}^{0}$ was introduced (but not so named) in [ABFR91], where it was shown that $\textsf{MAC}^{0}\subsetneq\textsf{TC}^{0}$. This class was subsequently given the name $\textsf{MAC}^{0}$ in [JKS02].
@@ 3.1 Additional preliminaries for groups with Abelian normal Hall subgroup
A $\mathbb{Z}H$-module is an abelian group $N$ together with an action of $H$ on $N$, given by a group homomorphism $\theta\colon H\to\operatorname{Aut}(N)$. We refer to these as “$H$-modules.” A submodule of an $H$-module NN is a subgroup $N^{\prime}\leq N$ such that the action of $H$ on $N^{\prime}$ sends $N^{\prime}$ into itself, and thus the restriction of the action of $H$ to $N^{\prime}$ gives $N^{\prime}$ the structure of an $H$-module compatible with that on $N$. Given a subset $S\subseteq N$, the smallest $H$-submodule containing $S$ is denoted $\langle S\rangle_{H}$, and is referred to as the $H$-submodule generated by $S$. An $H$-module generated by a single element is called cyclic. Note that a cyclic $H$-module NN need not be a cyclic Abelian group.
Two $H$-modules $N,N^{\prime}$ are isomorphic (as $H$-modules), denoted $N\cong_{H}N^{\prime}$, if there is a group isomorphism $\varphi\colon N\to N^{\prime}$ that is $H$-equivariant, in the sense that $\varphi(\theta(h)(n))=\theta^{\prime}(h)(\varphi(n))$ for all $h\in H,n\in N$. An $H$-module NN is decomposable if $N\cong_{H}N_{1}\oplus N_{2}$ where $N_{1},N_{2}$ are nonzero $H$-modules (and the direct sum can be thought of as a direct sum of Abelian groups); otherwise $N$ is indecomposable. An equivalent characterization of $N$ being decomposable is that there are nonzero $H$-submodules $N_{1},N_{2}$ such that $N=N_{1}\oplus N_{2}$ as Abelian groups (that is, $N$ is generated as a group by $N_{1}$ and $N_{2}$, and $N_{1}\cap N_{2}=0$). The Remak–Krull–Schmidt Theorem says that every $H$-module decomposes as a direct sum of indecomposable modules, and that the multiset of $H$-module isomorphism types of the indecomposable modules appearing is independent of the choice of decomposition, that is, it depends only on the $H$-module isomorphism type of $N$. We may thus write